OpenStudy (anonymous):

Find the line that is tangent to the curve y=x^3 at the point (1,1) The deravitive of the function is f'(x)=-3X^2

OpenStudy (anonymous):

Y-1=-3(x^2-1) y=-3x^2+4 This is what I am getting. The real answer is y=3x+2

OpenStudy (anonymous):

actually the derivative is \(f'(x)=3x^2\) and so at \(x=1\) the slope is \(f'(1)=3\)

OpenStudy (anonymous):

don't forget the derivative is not itself a slope, but rather a formula for a slope plug in the \(x\) value, in this case 1, and find the slope. the slope is a number, not a formula

OpenStudy (anonymous):

therefore you are asked to find the equation of a line with slope \(3\) through the point \((1,1)\) the point-slope formula will complete this but again i stress, \(f'(x)=3x^2\) is not itself the slope, it is the formula that gives it

OpenStudy (anonymous):

for example at \((2,8)\) the slope is \(f'(2)=3\times 2^2=12\) etc

OpenStudy (anonymous):

I think I understand. Thanks for fixing both my errors.

OpenStudy (anonymous):

yw